OpenStudy (twizttiez):

Heidi hikes two trails, the Big Trapezoid and the Little Trapezoid. The shorter base of the Big Trapezoid is 2 mi long, while the longer base is 8 mi. The longer base of the Little Trapezoid is 4 mi long. The trails are similar shapes. What is the length of the shorter base of the Little Trapezoid trail? mi

OpenStudy (anonymous):

Hi again.. XD also Since they're similar, their corresponding sides are proportional. So 8/4 = 2/x; solving, x = 1 mile https://answers.yahoo.com/question/index?qid=20150526133248AACvuHW this may help

OpenStudy (anonymous):

@TwiztTiez

OpenStudy (twizttiez):

Thanks!